Using first principle show that the derivative of COSX = -SINX


1
Expert's answer
2021-06-08T17:44:37-0400

"We \space will \space find \space derivative \\ \space of \space cos (x) with \space first \space principle.\\
First \space principle \space formula \space of \\ \space derivative \space such \space that...\\

f'(x)=\\\displaystyle \lim_{h\rightarrow 0}\;\;\;\frac{f(x+h)-(x)}{h}\\\\ =\displaystyle \lim_{h\rightarrow 0}\;\;\;\frac{cos(x+h)-cos(x)}{h}\\\\ =\displaystyle \lim_{h\rightarrow 0}\;\;\;\frac{cosx*cosh-sinx*sinh-cos(x)}{h}\\\\ =\displaystyle \lim_{h\rightarrow 0}\;\;\;\frac{cosx(cosh-1)-sinx*sinh}{h}\\\\ =\displaystyle \lim_{h\rightarrow 0}\;\;\;\frac{cosx(cosh-1)}{h}-\frac{sinx*sinh}{h}\\\\ =\displaystyle \lim_{h\rightarrow 0}\;\;\;cosx*\frac{(cosh-1)}{h}-sinx*\frac{sinh}{h}\\\\ =cosx*0-sinx*1\\\\ =0-sin(x)\\\\ =-sin(x)"